### v3

@@ Abstract
Keywords:
Out-of-Distribution Detection Global Pooling Layer Feature Activation Analysis
@@ 1 Introduction
In this study, we propose a novel prior, called Neural Activation Prior (NAP), for OOD detection. NAP characterizes our key observation that for channels before the global pooling layer in a fully trained neural network (illustrated to Figure 2(a)), a few neurons have a significantly high probability to show a larger response when activated by in-distribution (ID) samples compared to OOD samples. An intuitive explanation for NAP is that channels in a model fully trained on an ID dataset play a role in detecting certain patterns in the input samples from the ID dataset. When such patterns are detected, a few neurons can be activated [13], resulting in larger responses. These large responses usually occur when the input sample is ID data, but when the input is OOD data, such responses are rarely observed. This is because the pattern that the neuron focuses on is unlikely to be present in the OOD data. To verify our proposed prior, we employed DenseNet architecture [15] on CIFAR-10 [20] and Texture [5] datasets, analyzing mean and maximum within-channel activations before global pooling at the penultimate layer. Figure 1 clearly demonstrates that ID samples showing significantly higher maximal activation values than OOD samples at equal average activation levels. This is not unique to the $333^{\text{rd}}$ channel but is observed across most channels.
Figure 1: Activation distribution difference between ID data and OOD data inside channel. This illustrates the difference in within-channel activation distribution between CIFAR-10 [20] (ID data) and the Texture [5] (OOD data) datasets. The horizontal axis represents the average activation value of the channel, while the vertical axis shows the maximal activation value of the channel. Interestingly, at the same mean activation levels, the ID data (CIFAR-10) show significantly higher maximum activation values within the channel compared to the OOD data (Texture). It’s important to note that this pattern is not unique to the $333^{\text{rd}}$ channel but is observed across most channels.
@@
It is worth noting that our proposed prior is orthogonal to that used in current OOD detection methods. In the OOD deteciton field, as shown in Figure 2(a), existing methods [11, 23, 24, 33, 34, 8] mainly focus on the outputs and weights of the neural network after global pooling, and use them to design scoring functions for OOD detection. In contrast, the prior we proposed is focus within the channels of the penultimate layer before global pooling. Since the information carried in our prior can be easily lost during the global pooling process, this demonstrates that our NAP is essentially complementary to the priors used in current OOD detection studies. This property is verified in subsequent experiments Section 6 in this paper. To this end, we would like to emphasize that the contribution of this paper would lie in how much improvement we can achieve by integrating our method with existing approaches, rather than a direct comparison with existing methods in the field.
@@
Figure 2: Score Distribution Visualization Using DenseNet on CIFAR-10 (ID) and iSun (OOD) The integration of (a) Energy Score and (b) NAP Score through multiplication yields the (c) Energy $\times$ NAP Score, demonstrating superior differentiation between ID and OOD datasets. The effectiveness of this approach is attributed to the orthogonal nature of the proposed NAP relative to conventional OOD detection methods exemplified by the Energy Score. This illustrates that a simple multiplicative combination with NAP enhances detection capability. Importantly, the objective here is not to outperform the Energy Score per se but to highlight the synergistic potential of NAP as a complementary enhancement to the Energy Score and similar methods.
Experimental results show that our method achieves state-of-the-art performance on CIFAR-10 [20], CIFAR-100 [20] and ImageNet datasets. Specifically, our method significantly reduces the false positive rate by 40.06% (from 15.05 to 9.02) on the CIFAR-10 dataset [20], a reduction of 37.89% (from 41.40 to 25.71) on the CIFAR-100 dataset [20], and a reduction of 16.26% (from 35.66 to 29.86) on the ImageNet dataset. The large drop in FPR95 highlights the effectiveness of our approach in different environments. The above experimental results demonstrate the power of our proposed prior. We believe that these findings will provide inspiration for other researchers and thus promote progress in the field of OOD detection.
@@
  • We demonstrate the state-of-the-art performance of our approach through extensive experiments across various datasets, including a reduction in FPR95 by up to 40.06%. These results underscore the method’s operational efficiency, simplicity of deployment, and overall efficacy.
@@ 2.1 OOD Detection
The OOD detection community has explored a variety of techniques to underscore the distinctions between ID and OOD samples. These methods encompass classification-based [16, 23, 3, 7, 12], density-based [48, 1, 27, 47, 18, 29, 19, 31], and distance-based approaches [35, 22, 26, 4, 45, 37, 36, 25], with classification-based techniques generally outperforming the other types [42]. In classification-based methods, the basic work of OOD detection starts with a simple and effective baseline: using the Maximum Softmax Probability (MSP) [12] to measure the probability that a certain sample is an ID sample. On this basis, early approaches [23, 14, 24] focused on developing enhanced OOD indicators derived from neural network outputs. In addition, some researchers have proposed strategies involving OOD sample generation [21] and gradient-based [23] techniques. Among these, certain post-hoc methods [11, 23, 24, 33, 34, 8, 44] are notable for their simplicity and because they do not necessitate changes in the training process or objectives. This feature is particularly valuable for implementing OOD detection in real production environments, where the additional cost and complexity associated with retraining would be unacceptable.
@@ 3 Neural Activation Prior
Our NAP is based on the following observation for OOD detection: for a channel located before the global pooling layer in a fully trained neural network, the likelihood that a small number of its neurons activated with a stronger response to an ID sample is significantly higher compared to an OOD sample. For the behavior in other layers of the neural network, refer to the discussion around Figure 5 in Section 5.
@@
    $\text{Max}(\mathbf{A}_{j})=\max_{k,l}\mathbf{A}_{jkl},$
    where $\text{Max}(\mathbf{A}_{j})$ is the maximum value among all elements in the activation vector $\mathbf{A}_{j}$. Notice that $\mathbf{A}_{jkl}$ is processed by ReLU (as shown in Figure 2(a)), hence $\mathbf{A}_{jkl}\geq 0$.
@@
    $\text{Mean}(\mathbf{A}_{j})=\frac{1}{h\times w}\sum_{k=1}^{h}\sum_{l=1}^{w}\mathbf{A}_{jkl},$
@@ 4.2 Design of Scoring Function
[…]
@@
[…]
@@
[…]
@@ 5.1 Evaluation on CIFAR Benchmarks
Table 1: Comparison with competitive post-hoc OOD detection method on CIFAR benchmarks. All values are percentages and are averaged over 6 OOD test datasets. Except for our proposed method, all other results are taken from [8]. The best result is highlighted in bold. Methods marked with an asterisk (*) are employed with the Energy Score [24]. Note: A. = Area Under the ROC Curve; F. = False Positive Rate at 95% True Positive Rate. Methods include MSP [11], ODIN [23], Mahalanobis distance [22], Energy [24], ReAct [33], DICE [34], and ASH [8].
Method MSP ODIN Maha. Energy ReAct DICE ASH-B ASH-S NAP-E NAP-A NAP-R
CIFAR-10 F. $\downarrow$ 48.73 24.57 31.42 26.55 26.45 20.83 20.23 15.05 9.02 11.14 9.18
A. $\uparrow$ 92.46 93.71 89.15 94.57 94.67 95.24 96.02 96.91 98.15 97.48 98.02
CIFAR-100 F. $\downarrow$ 80.13 58.14 55.37 68.45 62.27 49.72 48.73 41.40 32.61 35.40 25.71
A. $\uparrow$ 74.36 84.49 82.73 81.19 84.47 87.23 88.04 90.02 92.84 91.21 93.18
[…]
Comparison. For comparison, we adopted recent post-hoc OOD detection methods: MSP [11], ODIN [23], Mahalanobis distance [22], Energy [24], ReAct [33], DICE [34], and ASH [8].
Experimental Results. Table 1 shows the comparison of NAP with other post hoc OOD detection methods on the CIFAR-10 and CIFAR-100 benchmarks. As shown in the table, our method significantly outperforms all other methods on the CIFAR-10 and CIFAR-100 datasets, achieving state-of-the-art performance. In CIFAR-10, NAP-E reduces FPR95 by 40.06% (from 15.05 to 9.02) compared to the previous best method. In CIFAR-100, NAP-R reduces FPR95 by 37.89% (from 41.40 to 25.71) compared to the previous best method. What is impressive is that the implementation of NAP-E is extremely simple and does not cause any degradation in ID data classification accuracy. This enables NAP-E to achieve the optimal pareto front among existing methods, where relevant results can be seen in Section 5.4. The Above results demonstrate the effectiveness of our method and the power of NAP in OOD detection.
Table 2: OOD detection results on ImageNet-1k [6]. All values are percentages. Except for our proposed method, all other results are taken from [8]. Methods marked with an asterisk (*) are employed with the Energy Score [24].
@@
iNat. [38] SUN [40] Places [46] Textures [5]
FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC FPR95 AUROC
@@
ODIN [23] 55.39 87.62 54.07 85.88 57.36 84.71 49.96 85.03 54.20 85.81
Maha. [22] 62.11 81.00 47.82 86.33 52.09 83.63 92.38 33.06 63.60 71.01
Energy [24] 59.50 88.91 62.65 84.50 69.37 81.19 58.05 85.03 62.39 84.91
ReAct* [33] 42.40 91.53 47.69 88.16 51.56 86.64 38.42 91.53 45.02 89.47
DICE* [34] 43.09 90.83 38.69 90.46 53.11 85.81 32.80 91.30 41.92 89.60
ASH-B* [8] 31.46 94.28 38.45 91.61 51.80 87.56 20.92 95.07 35.66 92.13
ASH-S* [8] 39.10 91.94 43.62 90.02 58.84 84.73 13.12 97.10 38.67 90.95
NAP-E 29.90 94.47 39.69 90.46 55.17 85.15 11.74 97.28 34.12 91.84
NAP-A 26.26 95.10 32.89 92.77 48.69 87.92 11.60 97.32 29.86 93.28
@@ 5.2 Evaluation on ImageNet
Implementation Details. In real-world applications, models are confronted with high-resolution images spanning a diverse range of scenes and features. Evaluations on large-scale datasets can provide insights into the performance of models in practical deployments. Thus, in line with recent research [33, 34, 8], we conduct experiments with NAP on the expansive ImageNet-1k [6] dataset in this study. Four dataset subsets, with all overlapping categories with ImageNet-1k eliminated, were employed as OOD benchmarks. These OOD datasets comprise Textures [5], Places365 [46], iNaturalist [38], and SUN [40]. We used MobileNetV2 [32] architecture, which pre-trained on ImageNet-1k [6]. The architecture and parameters remain unchanged during the OOD detection stage. Parallel to the experiments conducted on the CIFAR [20] benchmarks, we experimented with three distinct implementations. For parameter settings, $w$ was set to $0.6$ for NAP-E, $0.8$ for NAP-A, and $0.8$ for NAP-R.
Experimental Results. Table 2 shows the comparison results of NAP with other post-hoc OOD detection methods on the ImageNet-1k [6] benchmark. As shown in the table, our method outperforms all other methods on the ImageNet-1k dataset [6], achieving state-of-the-art performance. We have observed that our method brings significant performance improvements on the iNaturalist and Textures datasets, which is particularly notable. Intuitively, most of the pictures in the Textures [5] dataset are of a texture nature, so there is a small probability of obtaining a large response value. And while samples in iNaturalist have a relatively simple background, and the animals and plants in the foreground have certain semantic differences from the samples in ImageNet-1k, therefor, a sample in iNaturalist [38] is not likely to trigger large response values. The reason why existing methods do not perform well on this dataset may be due to their focus on activation values after global pooling. We speculate that although texture pictures cannot cause large activation values, they can excite small-amplitude noise on the entire feature map. Since large response values tend to only appear in a small area on the feature map, after global pooling, ID samples and texture class samples are likely to obtain similar average activation values. This may compromise the separability of the two types of samples, causing existing work to perform poorly on such samples.

### v4

@@ 1 Introduction
In this study, we propose a novel prior, called Neural Activation Prior (NAP), for OOD detection. NAP characterizes our key observation that for channels before the global pooling layer in a fully trained neural network (illustrated to Figure 3(a)), a few neurons have a significantly high probability to show a larger response when activated by in-distribution (ID) samples compared to OOD samples. An intuitive explanation for NAP is that channels in a model fully trained on an ID dataset play a role in detecting certain patterns in the input samples from the ID dataset. When such patterns are detected, a few neurons can be activated [16], resulting in larger responses. These large responses usually occur when the input sample is ID data, but when the input is OOD data, such responses are rarely observed. This is because the pattern that the neuron focuses on is unlikely to be present in the OOD data. To verify our proposed prior, we employed DenseNet architecture [18] on CIFAR-10 [23] and Texture [5] datasets, analyzing mean and maximum within-channel activations before global pooling at the penultimate layer. Figure 1 clearly demonstrates that ID samples exhibit significantly higher maximal activation values than OOD samples at equal average activation levels. Consequently, a series of methods, such as those in [14, 40, 27, 26, 25, 39, 8, 2, 38], based on pooled activation values are unable to effectively distinguish these OOD samples, given the non-discriminative nature of their average activation values.
(a) Activation Maps of ID and OOD samples.
(b) Activation distribution difference between ID data and OOD data.
Figure 1: Global pooling leads to the neglect of distributional characteristics of activation values within channels, thereby making it difficult to distinguish between ID and OOD samples. (a) illustrates the activation value maps in the penultimate layer for in-distribution (ID) samples (left) and out-of-distribution (OOD) samples (right) within the same channel. In this layer, different channels typically focus on different semantic features. When a specific feature appears in an image, such as the central region in the left image, that location will have very high activation values. Although OOD samples lack these specific features, the model’s unfamiliarity with OOD data leads to unpredictable activations, possibly causing weak noise activations (right image). This is detailed in [38]. Existing methods often rely on pooled activation values for OOD detection. Thus, in (a), both ID and OOD samples have a mean channel activation of 1.15, making them indistinguishable by existing methods. However, the NAP score proposed in this paper can effectively distinguish between them. More examples like (a) are provided in the Appendix J. (b) shows the activation distribution differences between ID and OOD data within the channel. The horizontal axis represents the average activation value, while the vertical axis represents the maximum activation value. Interestingly, at the same average activation level, ID data (CIFAR-10) shows significantly higher maximum activation values than OOD data (textures). This pattern is not unique to the $333^{\text{rd}}$ channel but is observable in most channels, corroborating the phenomenon in (a).
@@
It is worth noting that our proposed prior is orthogonal to that used in current OOD detection methods. In the OOD deteciton field, as shown in Figure 3(a), existing methods [14, 40, 27, 26, 25, 39, 8, 2, 38] mainly focus on the outputs and weights of the neural network after global pooling, and use them to design scoring functions for OOD detection. In contrast, the prior we proposed is focus within the channels of the penultimate layer before global pooling. Since the information carried in our prior can be easily lost during the global pooling process, this demonstrates that our NAP is essentially complementary to the priors used in current OOD detection studies. To this end, we would like to emphasize that the contribution of this paper would lie in how much improvement we can achieve by integrating our method with existing approaches, rather than a direct comparison with existing methods in the field.
@@
Figure 2: Score distribution visualization using DenseNet on CIFAR-10 (ID) and iSun (OOD) The integration of (a) Energy Score and (b) NAP Score through multiplication yields the (c) Energy $\times$ NAP Score, demonstrating superior differentiation between ID and OOD datasets. The effectiveness of this approach is attributed to the orthogonal nature of the proposed NAP relative to conventional OOD detection methods exemplified by the Energy Score. This illustrates that a simple multiplicative combination with NAP enhances detection capability. Importantly, the objective is not merely to surpass the performance of the Energy Score itself, but to underscore the synergistic potential of NAP as a complementary enhancement to the Energy Score and similar method.
Experimental results show that our method achieves state-of-the-art performance on CIFAR-10 [23], CIFAR-100 [23] and ImageNet datasets. Specifically, our method significantly reduces the false positive rate by 48.23% (from 15.05 to 7.79) on the CIFAR-10 dataset [23], a reduction of 37.89% (from 41.40 to 25.71) on the CIFAR-100 dataset [23], and a reduction of 16.26% (from 35.66 to 29.86) on the ImageNet dataset. The large drop in FPR95 highlights the effectiveness of our approach in different environments. The above experimental results demonstrate the power of our proposed prior. We believe that these findings will provide inspiration for other researchers and thus promote progress in the field of OOD detection.
@@
  • We demonstrate the state-of-the-art performance of our approach through extensive experiments across various datasets, including a reduction in FPR95 by up to 48.23%. These results underscore the method’s operational efficiency, simplicity of deployment, and overall efficacy.
@@ 2.1 OOD detection
The OOD detection community has explored a variety of techniques to underscore the distinctions between ID and OOD samples. These methods encompass classification-based [19, 26, 3, 7, 15, 36, 42, 10], density-based [55, 1, 30, 54, 21, 32, 22, 34], and distance-based approaches [41, 25, 29, 4, 52, 44, 43, 28, 40], with classification-based techniques generally outperforming the other types [49]. In classification-based methods, the basic work of OOD detection starts with a simple and effective baseline: using the Maximum Softmax Probability (MSP) [15] to measure the probability that a certain sample is an ID sample. On this basis, early approaches [26, 17, 27] focused on developing enhanced OOD indicators derived from neural network outputs. In addition, some researchers have proposed strategies involving OOD sample generation [24, 10] and gradient-based [26] techniques. Among these, certain post-hoc methods [14, 26, 27, 38, 39, 8, 51, 10] are notable for their simplicity and because they do not necessitate changes in the training process or objectives. This feature is particularly valuable for implementing OOD detection in real production environments, where the additional cost and complexity associated with retraining would be unacceptable.
@@ 3 Neural activation prior
Our NAP is based on the following observation for OOD detection: for a channel located before the global pooling layer in a fully trained neural network, the likelihood that a small number of its neurons activated with a stronger response to an ID sample is significantly higher compared to an OOD sample. For the behavior in other layers of the neural network, refer to the discussion around Figure 4 in Section 5.4.
@@
    $\text{Max}(\mathbf{A}_{j})=\max_{k,l}\mathbf{A}_{jkl},$ (1)
    where $\text{Max}(\mathbf{A}_{j})$ is the maximum value among all elements in the activation vector $\mathbf{A}_{j}$. Here, $k$ and $l$ index the spatial dimensions (height and width) of the activation map, respectively. Notice that $\mathbf{A}_{jkl}$ is processed by ReLU (as shown in Figure 3(a)), hence $\mathbf{A}_{jkl}\geq 0$.
@@
    $\text{Mean}(\mathbf{A}_{j})=\frac{1}{h\times w}\sum_{k=1}^{h}\sum_{l=1}^{w}\mathbf{A}_{jkl},$ (2)
@@ 4.2 Design of scoring function
$S_{\textit{NAP}}(x;f)=\frac{1}{C}\sum_{j=1}^{C}\left(\frac{\text{Max}(\mathbf{A}_{j})}{\text{Mean}(\mathbf{A}_{j})+\epsilon}\right)^{2},$ (3)
@@
OOD detection. The usage of scoring function $S_{\textit{NAP}}(x;f)$ in this paper is similar to that of the energy score $-E(x;f)$. The energy scoring function $E(x;f)$ converts the logits output of the classification network $f$ into a scalar $E(x;f)=-\log\sum_{i=1}^{K}e^{f_{i}(x)}$, where $f_{i}(x)$ is the logits output of category $i$. In OOD detection, the score employed for OOD detection is the negative energy score, $-E(x;f)$. Therefore, ID data is given a higher score, while OOD data is assigned a lower score.
@@
$S_{\textit{NAP-E}}(x;f,w)=-E(x;f)^{w}\cdot S_{\textit{NAP}}(x;f)^{1-w}.$ (4)
And when using $S_{\textit{NAP}}(x;f)$ to enhance other energy score based OOD methods, we simply replace the function $f$ in the above formula with the specific function of the corresponding method, such as $f^{\textit{DICE}}$, $f^{\textit{ReAct}}$, $f^{\textit{ASH}}$, etc.
How to find a optimal parameter $w$? When combining NAP with different OOD detection methods, the optimal weight parameter $w$ varies. To obtain the optimal parameter, we utilized a set of data transformation techniques (such as Gaussian noise, glass blur, motion blur, etc., more details in the Appendix H) to generate a corrupted dataset based on the ID dataset, serving as pseudo OOD data. For the choice of transformation types, we referred to [13]. Utilizing this set of OOD data, we employed a binary search method to find the optimal $w$. Through experimentation with various datasets and methods, we found that this search approach quickly identifies the optimal $w$, which generalizes well to real OOD datasets. Refer to Appendix H for more details about this process.
@@ 5 Experiments
In this section, we conduct experiments on various real-world datasets. In our experiments, we use NAP-[initial] to denote the combination of NAP with another method, where [initial] represents the initial letter of the method’s name (e.g., NAP-A for the combination with ASH). Specifically, we combine NAP with a series of common OOD detection methods, including ASH, DICE, Energy, KNN, MSP, and ReAct, denoted as NAP-A, NAP-D, NAP-E, NAP-K, NAP-M, and NAP-R, respectively. CIFAR-10, CIFAR-100, and ImageNet are used as ID datasets. For each combination of NAP with other methods, we use the approach described in Section 4.2 to determine the optimal combination parameter $w$. Detailed optimal $w$ values for different experimental setups can be found in Appendix H. It’s noted that the value of $\epsilon$ in our scoring functions is consistently set to $1.0$ for numerical stability. All experiments were conducted on an NVIDIA GeForce RTX 3090 GPU.
@@ 5.1 Evaluation on CIFAR benchmarks
Table 1: Comparison with competitive post-hoc OOD detection method on CIFAR benchmarks. All values are percentages and are averaged over 6 OOD test datasets. Note: A. = Area Under the ROC Curve; F. = False Positive Rate at 95% True Positive Rate. Methods include ASH [8], DICE [39], Energy [27], KNN [40], MSP [14], and ReAct [38].
Method ASH NAP-A DICE NAP-D Energy NAP-E KNN NAP-K MSP NAP-M ReAct NAP-R
F. $\downarrow$ 15.05 11.14 20.83 11.66 26.55 9.02 16.12 7.79 48.69 19.09 26.45 9.18
CIFAR-10 A. $\uparrow$ 96.91 97.48 95.24 97.47 94.67 98.15 96.79 98.38 92.52 95.11 94.67 98.02
F. $\downarrow$ 41.40 35.40 49.72 32.34 68.45 32.61 44.91 33.63 80.13 48.20 62.27 25.71
CIFAR-100 A. $\uparrow$ 90.02 91.21 87.23 92.23 81.19 92.84 86.58 91.54 74.36 88.45 84.47 93.18
Experimental results. Table 1 presents the comparison of NAP combined with other post hoc OOD detection methods on the CIFAR-10 and CIFAR-100 benchmarks. As shown in the table, our approach significantly enhances the performance of all methods on both CIFAR-10 and CIFAR-100 datasets. Notably, the maximum reductions in FPR95 on CIFAR-10 and CIFAR-100 are 66.03% (NAP-E) and 58.71% (NAP-R), respectively. Note that the table showcases the average performance across six OOD datasets; for complete performance details, please refer to Appendix G.
Table 2: OOD detection results on ImageNet-1k [6]. All values are percentages. Methods include ASH [8], DICE [39], Energy [27], KNN [40], MSP [14], and ReAct [38].
@@
iNaturalist [45] SUN [47] Places [53] Textures [5]
FPR95 $\downarrow$ AUROC $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$ FPR95 $\downarrow$ AUROC $\uparrow$
ASH [8] 31.46 94.28 38.45 91.61 51.80 87.56 20.92 95.07 35.66 92.13
NAP-A 26.26 95.10 32.89 92.77 48.69 87.92 11.60 97.32 29.86 93.28
DICE [39] 43.09 90.83 38.69 90.46 53.11 85.81 32.80 91.30 41.92 89.60
NAP-D 27.48 94.13 36.14 90.66 51.84 85.03 9.02 97.92 31.12 91.94
Energy [27] 59.50 88.91 62.65 84.50 69.37 81.19 58.05 85.03 62.39 84.91
NAP-E 29.90 94.47 39.69 90.46 55.17 85.15 11.74 97.28 34.12 91.84
KNN [40] 85.91 72.67 90.49 65.39 93.18 60.08 14.08 96.98 70.92 73.78
NAP-K 38.23 89.80 56.55 80.01 70.89 71.35 7.02 98.39 43.17 84.89
@@
NAP-M 35.47 92.53 51.19 86.51 63.77 80.61 15.14 97.09 41.39 89.19
ReAct [38] 42.40 91.53 47.69 88.16 51.56 86.64 38.42 91.53 45.02 89.47
@@ 5.2 Evaluation on ImageNet
Implementation details. In real-world applications, models are confronted with high-resolution images spanning a diverse range of scenes and features. Evaluations on large-scale datasets can provide insights into the performance of models in practical deployments. Thus, in line with recent research [38, 39, 8], we conduct experiments with NAP on the expansive ImageNet-1k [6] dataset in this study. Four dataset subsets, with all overlapping categories with ImageNet-1k eliminated, were employed as OOD benchmarks. These OOD datasets comprise Textures [5], Places365 [53], iNaturalist [45], and SUN [47]. We used MobileNetV2 [35] architecture, which pre-trained on ImageNet-1k [6]. The architecture and parameters remain unchanged during the OOD detection stage.
Experimental results. Table 2 shows the comparison results of NAP with other post-hoc OOD detection methods on the ImageNet-1k [6] benchmark. As shown in the table, our approach significantly enhances the performance of all methods on the ImageNet-1k dataset [6]. We have observed that our method brings significant performance improvements on the iNaturalist and Textures datasets, which is particularly notable. Intuitively, most of the pictures in the Textures [5] dataset are of a texture nature, so there is a small probability of obtaining a large response value. And while samples in iNaturalist have a relatively simple background, and the animals and plants in the foreground have certain semantic differences from the samples in ImageNet-1k, therefor, a sample in iNaturalist [45] is not likely to trigger large response values. The reason why existing methods do not perform well on this dataset may be due to their focus on activation values after global pooling. We speculate that although texture pictures cannot cause large activation values, they can excite small-amplitude noise on the entire feature map. Since large response values tend to only appear in a small area on the feature map, after global pooling, ID samples and texture class samples are likely to obtain similar average activation values. This may compromise the separability of the two types of samples, causing existing work to perform poorly on such samples.